# Sparse Covariance Neural Networks

### v1

@@ C.1 Preliminaries
Term 3. This term depends on the second-order error $((\bar{\lambda}_{i}-\lambda_{i})\mathbf{I}_{N}-\mathbf{E})(\mathbf{\bar{v}}_{i}-\mathbf{v}_{i})$. From Weyl’s theorem (Golub & van Loan, 2013, Theorem 8.1.6), we have that $\|(\bar{\lambda}_{i}-\lambda_{i})\mathbf{I}_{N}-\mathbf{E}\|\leq 2\|\mathbf{E}\|$. From equation 18, we have that $\|\mathbf{\bar{v}}_{i}-\mathbf{v}_{i}\|=\mathcal{O}(\|\mathbf{E}\|)$. Therefore, the product of the two is $\|(\bar{\lambda}_{i}-\lambda_{i})\mathbf{I}_{N}-\mathbf{E}\|\|\mathbf{\bar{v}}_{i}-\mathbf{v}_{i}\|\approx\mathcal{O}(\|\mathbf{E}\|^{2})$, which is dominated by the other two terms in equation 10 and equation 11 under small perturbation assumption.
We complete the proof by merging the terms in equation 24 and equation 16.
C.2 Proof of Thm. 2
From (Bickel & Levina, 2008b, Theorem 1), given a true covariance $\mathbf{C}$ and a hard-thresholded sample covariance $\mathbf{\bar{C}}$, it holds with high probability that
$\|\mathbf{\bar{C}}-\mathbf{C}\|=\|\mathbf{E}\|\leq c_{0}\sqrt{\frac{\log N}{t}}.$ (25)
By replacing $\|\mathbf{E}\|$ in equation 9, the claim follows. ∎
C.3 Proof of Thm. 3
From (Deshpande & Montanari, 2016, Theorem 1), given a true covariance $\mathbf{C}$ and a soft-thresholded sample covariance $\mathbf{\bar{C}}$, it holds with high probability that
@@
for a generic constant $C$. By replacing $\|\mathbf{E}\|$ in equation 9, the claim follows. ∎
Appendix D Proof of Proposition 1
@@
where we used triangle inequality, the fact that $\|\mathbf{x}\|\leq 1$ and the fact that $\|(\mathbf{v}_{i}-\mathbf{\bar{v}}_{i})^{\mathsf{T}}\|=\|\mathbf{v}_{i}-\mathbf{\bar{v}}_{i}\|$. Using the sin-theta theorem (Stewart & guang Sun, 1990, Theorem V.3.6) (cf. equation 18), we have that $\max_{i}\left\|\mathbf{v}_{i}-\mathbf{\bar{v}}_{i}\right\|\leq\sqrt{2}\|\mathbf{E}\|/\min_{i}|\lambda_{i}-\lambda_{i+1}|$ where $\|\mathbf{E}\|=\|\mathbf{C}-\mathbf{\bar{C}}\|$. Now, we use the expression for $\|\mathbf{E}\|$ with hard thresholding in (Bickel & Levina, 2008b, Theorem 1), i.e.,
@@
Appendix E Generalized Covariance Filter Frequency Response
We provide here additional details related to the derivation of the generalized covariance filter frequency response in Def. 6 following (Gao et al., 2021a). Consider a covariance filter operating on a sequence of $K$ sparsified covariances $\mathbf{\tilde{C}}_{1}\dots\mathbf{\tilde{C}}_{K}$ with eigendecomposition $\mathbf{\tilde{C}}_{k}=\mathbf{\tilde{V}}_{k}\mathbf{\tilde{\mathbf{\Lambda}}}_{k}\mathbf{\tilde{V}}_{k}^{\mathsf{T}}$ (where $\mathbf{\tilde{V}}_{k}=[\mathbf{\tilde{v}}_{k1},\dots,\mathbf{\tilde{v}}_{kN}]$ and $\mathbf{\tilde{\mathbf{\Lambda}}}_{k}=\text{diag}(\tilde{\lambda}_{k1},\dots,\tilde{\lambda}_{kN})$) according to Def. 5. For matrix $\mathbf{\tilde{C}}_{1}$, we can express a signal $\mathbf{x}$ as $\mathbf{x}=\sum_{i_{1}=1}^{N}\hat{x}_{1i_{1}}\mathbf{\tilde{v}}_{1i_{1}}$, where $\mathbf{\hat{x}}_{1}=[\hat{x}_{11},\dots,\hat{x}_{1N}]^{\mathsf{T}}$ is the covariance Fourier transform of $\mathbf{x}$ w.r.t. $\mathbf{\tilde{C}}_{1}$. By performing a graph signal shift, we obtain
@@ Appendix E Generalized Covariance Filter Frequency Response
When performing a second shift, i.e., $\mathbf{x}^{(2)}=\mathbf{\tilde{C}}_{2}\mathbf{\tilde{C}}_{1}\mathbf{x}$, we decompose each eigenvector $\mathbf{\tilde{v}}_{1i_{1}}$ by taking its covariance Fourier transform w.r.t. $\mathbf{\tilde{C}}_{2}$, i.e., we get $\mathbf{\tilde{v}}_{1i_{1}}=\sum_{i_{2}=1}^{N}\hat{x}_{2i_{1}i_{2}}\mathbf{\tilde{v}}_{2i_{2}}$ where $\mathbf{\tilde{x}}_{2i_{1}}=[\hat{x}_{2i_{1}1},\dots,\hat{x}_{2i_{1}N}]$ is the covariance Fourier transform of $\mathbf{\hat{v}}_{1i_{1}}$ over $\mathbf{\tilde{C}}_{2}$. Performing this operation on all eigenvectors $\mathbf{\tilde{v}}_{11},\dots,\mathbf{\tilde{v}}_{1N}$ and using equation 29, we can write
$\mathbf{x}^{(2)}=\mathbf{\tilde{C}}_{2}\sum_{i_{1}=1}^{N}\hat{x}_{1i_{1}}\tilde{\lambda}_{1i_{1}}\mathbf{\tilde{v}}_{1i_{1}}=\sum_{i_{2}=1}^{N}\sum_{i_{1}=1}^{N}\hat{x}_{2i_{1}i_{2}}\hat{x}_{1i_{1}}\tilde{\lambda}_{2i_{2}}\tilde{\lambda}_{1i_{1}}\mathbf{\tilde{v}}_{1i_{1}}.$ (30)
@@
and by aggregating the $K+1$ shifted signals according to Def. 5 we obtain
@@
From equation 32 we see that this representation involves all eigenvalues $\mathbf{\tilde{\mathbf{\Lambda}}}_{K},\dots,\mathbf{\tilde{\mathbf{\Lambda}}}_{1}$ and eigenvectors $\mathbf{\tilde{V}}_{K},\dots,\mathbf{\tilde{V}}_{1}$ of the sequence of covariance realizations. Therefore, we can consider the coefficients $\{\hat{x}_{1i_{1}}\}_{i_{1}=1}^{N}$ and $\{\hat{x}_{2i_{j}i_{j+1}}\}_{j=1}^{K-1}$ as the generalized covariance Fourier transform of signal $\mathbf{x}$ on the sequence of sparsified covariances $\mathbf{\tilde{V}}_{K},\dots,\mathbf{\tilde{V}}_{1}$. This supports the definition of a generalized frequency response for a covariance filter over random sparsified covariances in Def. 6.
Appendix F Proof of Thm. 4
@@
$\|\mathbf{H}(\mathbf{C})\mathbf{x}-\mathbf{H}(\mathbf{\tilde{C}})\mathbf{x}\|^{2}=\|\mathbf{H}(\mathbf{C})\mathbf{x}-\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}+\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}-\mathbf{H}(\mathbf{\tilde{C}})\mathbf{x}\|^{2}\leq$ (33)
$\|\mathbf{H}(\mathbf{C})\mathbf{x}-\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}\|^{2}+\|\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}-\mathbf{H}(\mathbf{\tilde{C}})\mathbf{x}\|^{2}=\alpha+\beta$ (34)
@@
F.1 Covariance uncertainty error
The effect of covariance estimation errors on the stability of VNN is deterministic and, consequently, can be bounded by the square of the bound in Thm. 1:
$\|\mathbf{H}(\mathbf{C})\mathbf{x}-\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}\|^{2}\leq\beta^{2}=\frac{P^{2}}{t}\mathcal{O}\left(N+\frac{\|\mathbf{C}\|^{2}\log(Nt)}{\nu^{2}t^{2}}\right)$ (35)
where $t$ is the number of data samples used to estimate the covariance matrix.
F.2 Covariance sparsification error
@@
Consider a random covariance matrix $\mathbf{\tilde{C}}_{r}=\mathbf{\hat{C}}+\mathbf{E}_{r}$, where $\mathbf{\tilde{C}}_{r}$ is a copy of the sample covariance matrix $\mathbf{\hat{C}}$ whose elements $\hat{c}_{ij}$ are dropped independently with probability $1-p_{ij}$ and $\mathbf{E}_{r}$ is its distance from $\mathbf{\hat{C}}$. Under the conditions of Def. 4, i.e., edge perturbations are undirected (i.e., $\mathbf{E}_{r}$ is symmetric) and no perturbations occur on the diagonal (i.e., $\mathbf{E}_{r}$ has zeros on the diagonal), it holds that
@@
Each entry of the random error matrix $\mathbf{E}_{r}$ can be represented as $[\mathbf{E}_{r}]_{ij}=-\delta_{ij}c_{ij}$, where $\delta_{ij}$ is a Bernoulli variable that is one with probability $1-p_{ij}$ and zero with probability $p_{ij}$. By performing the matrix multiplication, we can express each element of $\mathbf{E}_{r}^{2}$ and, consequently, its expectation, as
@@
$\mathsf{trace}(\mathbb{E}[\mathbf{E}_{r}^{2}])=\sum_{i=1}^{N}\sum_{n=1}^{N}\hat{c}_{in}\hat{c}_{ni}\mathbb{E}[\delta_{in}\delta_{ni}]=\hat{c}_{in}^{2}(1-p_{in}).$ (39)
@@
$\mathbb{E}\left[\sum_{r=1}^{K}\mathsf{trace}\left(\sum_{k=r}^{K}\sum_{l=r}^{K}h_{k}h_{l}\mathbf{E}_{r}\mathbf{\hat{C}}^{k+l-2r}\mathbf{E}_{r}\mathbf{\hat{C}}^{r-1}\mathbf{x}\mathbf{x}^{\mathsf{T}}\mathbf{\hat{C}}^{r-1}\right)\right]\leq NP^{2}\|\mathbf{x}\|^{2}Q$ (40)
with $Q$ defined in equation 36.
@@
Following (Gao et al., 2021a, eq. (B.13)-(B.15)), we rewrite the term in equation 40 as
@@
which holds for any square matrix $\mathbf{A}$ and positive semi-definite matrix $\mathbf{B}$ (Wang et al., 1986). We note that $\mathbb{E}[\mathbf{E}_{r}^{2}]$ is positive semi-definite since $\mathbf{E}_{r}^{2}$ is the square of a symmetric matrix. Therefore, we use equation 42 to write
@@
From (Gao et al., 2021a, eq. (B.24)), we have that the term within the norm is bounded by the generalized Lipschitz coefficient of the filter, i.e.,
@@
Then, using Lemma 3, we have that
@@
Substituting these two terms in equation 44 leads to
@@
Similarly to Lemma 3, we can express each element of the expected value of the matrix product $\mathbf{E}_{r_{1}}\mathbf{E}_{r_{2}}$ as
@@
From Lemma 3 we have that
@@
$[\mathbb{E}[\mathbf{E}_{r_{1}}^{2}\mathbf{E}_{r_{2}}]]_{ij}=\sum_{m=1}^{N}\mathbb{E}[[\mathbf{E}_{r_{1}}^{2}]_{im}\hat{c}_{mj}\delta_{r_{2},mj}]=\sum_{m=1}^{N}\sum_{n=1}^{N}\hat{c}_{in}\hat{c}_{nm}\hat{c}_{mj}\mathbb{E}[\delta_{r_{1},in}\delta_{r_{1},nm}\delta_{r_{2},mj}].$ (53)
@@
$\mathbb{E}[\delta_{r_{1},in}\delta_{r_{1},nm}\delta_{r_{2},mj}]=\begin{cases}(1-p_{in})(1-p_{nm})(1-p_{mj})&\text{if }i\neq m\\ (1-p_{in})(1-p_{mj})&\text{if }i=m\end{cases}$ (54)
@@
$\mathsf{trace}(\mathbb{E}[\mathbf{E}_{r_{1}}^{2}\mathbf{E}_{r_{2}}])=\sum_{i=1}^{N}\sum_{m=1}^{N}\sum_{n=1}^{N}\hat{c}_{in}\hat{c}_{nm}\hat{c}_{mi}\mathbb{E}[\delta_{r_{1},in}\delta_{r_{1},nm}\delta_{r_{2},mi}].$ (55)
Therefore, the trace is a sum of terms of quadratic and cubic order in the probability value, i.e., $\mathcal{O}((1-p_{1})(1-p_{2}))$ or $\mathcal{O}((1-p_{1})(1-p_{2})(1-p_{3}))$, where $p_{1},p_{2},p_{3}$ are generic probabilities values and the quadratic terms dominate the behavior since $11-p<$ 1. So,
@@
Consider the sample covariance matrix $\mathbf{\hat{C}}$ and a random sparsified covariance $\mathbf{\tilde{C}}_{r}=\mathbf{\hat{C}}+\mathbf{E}_{r}$ as in Def. 4. We have that
@@
Similarly to Lemma 3, we can write the expected value of each element of $\mathbf{E}^{3}_{r}$ as
$[\mathbb{E}[\mathbf{E}^{3}_{r}]]_{ij}=\sum_{m=1}^{N}\mathbb{E}[[\mathbf{E}_{r}^{2}]_{im}\hat{c}_{mj}\delta_{mj}]=\sum_{m=1}^{N}\sum_{n=1}^{N}\hat{c}_{in}\hat{c}_{nm}\hat{c}_{mj}\mathbb{E}[\delta_{in}\delta_{nm}\delta_{mj}]$ (58)
@@
$\mathbb{E}[\delta_{in}\delta_{nm}\delta_{mi}]=\begin{cases}(1-p_{in})(1-p_{nm})(1-p_{mi})&\text{if }i\neq m\land i\neq n\\ (1-p_{in})(1-p_{mi})&\text{if }i=m\oplus i=n\\ (1-p_{in})&\text{if }i=m=n\end{cases}$ (60)
where $\oplus$ denotes the xor operator. Now we note that, according to Def. 4, the matrix $\mathbf{E}_{r}$ has zeros on the diagonal or, equivalently, $1-p_{ii}=0\quad\forall i$. As a consequence, the terms where $i=m=n$ in equation 59 are all zeros and the trace only contains terms of quadratic and cubic order in the probability value, i.e., $\mathcal{O}((1-p_{1})(1-p_{2}))$ or $\mathcal{O}((1-p_{1})(1-p_{2})(1-p_{3}))$, where $p_{1},p_{2},p_{3}$ are generic probabilities values and the quadratic terms dominate the behavior since $11-p<$ 1. Therefore,
@@
$\mathbb{E}[\|\mathbf{u}-\mathbf{\tilde{u}}\|^{2}]=\mathbb{E}[\mathsf{trace}(\mathbf{u}^{\mathsf{T}}\mathbf{u}+\mathbf{\tilde{u}}^{\mathsf{T}}\mathbf{\tilde{u}}-2\mathbf{u}^{\mathsf{T}}\mathbf{\tilde{u}})]=$ (62)
$\mathbb{E}[\mathsf{trace}(\mathbf{\tilde{u}}^{\mathsf{T}}\mathbf{\tilde{u}}-\mathbf{u}^{\mathsf{T}}\mathbf{u})]+2\mathbb{E}[\mathsf{trace}(\mathbf{u}^{\mathsf{T}}\mathbf{u}-\mathbf{u}^{\mathsf{T}}\mathbf{\tilde{u}})]$ (63)
@@
We represent a random covariance matrix as $\mathbf{\tilde{C}}_{r}=\mathbf{\hat{C}}+\mathbf{E}_{r}$, where $\mathbf{E}_{r}$ is a random matrix that contains the deviation from the true covariance. Following (Gao et al., 2021a, eq. (B.2)-(B.6)), we express the first term in equation 63 as
$\mathbb{E}[\mathsf{trace}(\mathbf{\tilde{u}}^{\mathsf{T}}\mathbf{\tilde{u}}-\mathbf{u}^{\mathsf{T}}\mathbf{u})]=-2\mathbb{E}[\mathsf{trace}(\mathbf{u}^{\mathsf{T}}\mathbf{u}-\mathbf{u}^{\mathsf{T}}\mathbf{\tilde{u}})]+$ (64)

### v2

@@ B-A Preliminaries
Term 3. This term depends on the second-order error $((\bar{\lambda}_{i}-\lambda_{i})\mathbf{I}_{N}-\mathbf{E})(\mathbf{\bar{v}}_{i}-\mathbf{v}_{i})$. From Weyl’s theorem [74, Theorem 8.1.6], we have that $\|(\bar{\lambda}_{i}-\lambda_{i})\mathbf{I}_{N}-\mathbf{E}\|\leq 2\|\mathbf{E}\|$. From (8), we have that $\|\mathbf{\bar{v}}_{i}-\mathbf{v}_{i}\|=\mathcal{O}(\|\mathbf{E}\|)$. Therefore, the product of the two is $\|(\bar{\lambda}_{i}-\lambda_{i})\mathbf{I}_{N}-\mathbf{E}\|\|\mathbf{\bar{v}}_{i}-\mathbf{v}_{i}\|\approx\mathcal{O}(\|\mathbf{E}\|^{2})$, which is dominated by the other two terms in (12) and (13) under small perturbation assumption.
We complete the proof by merging the terms in (23),(18).
B-B Proof of Thm.˜3
From [25, Theorem 1], given a true covariance $\mathbf{C}$, a hard-thresholded sample covariance $\mathbf{\bar{C}}$ and a large enough constant $C$, it holds with probability 1o(1)1-o(1) that
$\|\mathbf{\bar{C}}-\mathbf{C}\|=\|\mathbf{E}\|\leq Cc_{0}\sqrt{\frac{\log N}{t}}.$ (24)
By replacing $\|\mathbf{E}\|$ in (11), the claim follows. ∎
B-C Proof of Thm.˜4
From [18, Theorem 1], given a true covariance $\mathbf{C}$ and a soft-thresholded sample covariance $\mathbf{\bar{C}}$, it holds with probability 1o(1)1-o(1) that
@@
for a large enough constant $C$. By replacing $\|\mathbf{E}\|$ in (11), the claim follows. ∎
Appendix C Proof of Proposition˜1
@@
where we used triangle inequality, the fact that $\|\mathbf{x}\|\leq 1$ and the fact that $\|(\mathbf{v}_{i}-\mathbf{\bar{v}}_{i})^{\mathsf{T}}\|=\|\mathbf{v}_{i}-\mathbf{\bar{v}}_{i}\|$. Using the sin-theta theorem [72, Theorem V.3.6] (cf. (6)), we have that $\max_{i}\left\|\mathbf{v}_{i}-\mathbf{\bar{v}}_{i}\right\|\leq\sqrt{2}\|\mathbf{E}\|/\min_{j}|\lambda_{j}-\lambda_{j+1}|$ where $\|\mathbf{E}\|=\|\mathbf{C}-\mathbf{\bar{C}}\|$. Now, we use the expression for $\|\mathbf{E}\|$ with hard thresholding in [25, Theorem 1], i.e.,
@@
Appendix D Generalized Covariance Filter Frequency Response
We provide here additional details related to the derivation of the generalized covariance filter frequency response in Def.˜6 following [36]. Consider a covariance filter operating on a sequence of $K$ sparsified covariances $\mathbf{\tilde{C}}_{1}\dots\mathbf{\tilde{C}}_{K}$ with eigendecomposition $\mathbf{\tilde{C}}_{k}=\mathbf{\tilde{V}}_{k}\mathbf{\tilde{\mathbf{\Lambda}}}_{k}\mathbf{\tilde{V}}_{k}^{\mathsf{T}}$ (where $\mathbf{\tilde{V}}_{k}=[\mathbf{\tilde{v}}_{k1},\dots,\mathbf{\tilde{v}}_{kN}]$ and $\mathbf{\tilde{\mathbf{\Lambda}}}_{k}=\text{diag}(\tilde{\lambda}_{k1},\dots,\tilde{\lambda}_{kN})$) according to Def.˜5. For matrix $\mathbf{\tilde{C}}_{1}$, we can express a signal $\mathbf{x}$ as $\mathbf{x}=\sum_{i_{1}=1}^{N}\hat{x}_{1i_{1}}\mathbf{\tilde{v}}_{1i_{1}}$, where $\mathbf{\hat{x}}_{1}=[\hat{x}_{11},\dots,\hat{x}_{1N}]^{\mathsf{T}}$ is the covariance Fourier transform of $\mathbf{x}$ w.r.t. $\mathbf{\tilde{C}}_{1}$. By performing a graph signal shift, we obtain
@@ Appendix D Generalized Covariance Filter Frequency Response
When performing a second shift, i.e., $\mathbf{x}^{(2)}=\mathbf{\tilde{C}}_{2}\mathbf{\tilde{C}}_{1}\mathbf{x}$, we decompose each eigenvector $\mathbf{\tilde{v}}_{1i_{1}}$ by taking its covariance Fourier transform w.r.t. $\mathbf{\tilde{C}}_{2}$, i.e., we get $\mathbf{\tilde{v}}_{1i_{1}}=\sum_{i_{2}=1}^{N}\hat{x}_{2i_{1}i_{2}}\mathbf{\tilde{v}}_{2i_{2}}$ where $\mathbf{\tilde{x}}_{2i_{1}}=[\hat{x}_{2i_{1}1},\dots,\hat{x}_{2i_{1}N}]$ is the covariance Fourier transform of $\mathbf{\hat{v}}_{1i_{1}}$ over $\mathbf{\tilde{C}}_{2}$. Performing this operation on all eigenvectors $\mathbf{\tilde{v}}_{11},\dots,\mathbf{\tilde{v}}_{1N}$ and using (28), we can write
$\mathbf{x}^{(2)}=\mathbf{\tilde{C}}_{2}\sum_{i_{1}=1}^{N}\hat{x}_{1i_{1}}\tilde{\lambda}_{1i_{1}}\mathbf{\tilde{v}}_{1i_{1}}=\sum_{i_{2}=1}^{N}\sum_{i_{1}=1}^{N}\hat{x}_{2i_{1}i_{2}}\hat{x}_{1i_{1}}\tilde{\lambda}_{2i_{2}}\tilde{\lambda}_{1i_{1}}\mathbf{\tilde{v}}_{2i_{2}}.$ (29)
@@
and by aggregating the $K+1$ shifted signals according to Def.˜5 we obtain
@@
From (31) we see that this representation involves all eigenvalues $\mathbf{\tilde{\mathbf{\Lambda}}}_{K},\dots,\mathbf{\tilde{\mathbf{\Lambda}}}_{1}$ and eigenvectors $\mathbf{\tilde{V}}_{K},\dots,\mathbf{\tilde{V}}_{1}$ of the sequence of covariance realizations. Therefore, we can consider the coefficients $\{\hat{x}_{1i_{1}}\}_{i_{1}=1}^{N}$ and $\{\hat{x}_{2i_{j}i_{j+1}}\}_{j=1}^{K-1}$ as the generalized covariance Fourier transform of signal $\mathbf{x}$ on the sequence of sparsified covariances $\mathbf{\tilde{V}}_{K},\dots,\mathbf{\tilde{V}}_{1}$. This supports the definition of a generalized frequency response for a covariance filter over random sparsified covariances in Def.˜6.
Appendix E Proof of Thm.˜5
@@
$\|\mathbf{H}(\mathbf{C})\mathbf{x}-\mathbf{H}(\mathbf{\tilde{C}})\mathbf{x}\|^{2}=$ (32)
$\|\mathbf{H}(\mathbf{C})\mathbf{x}-\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}+\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}-\mathbf{H}(\mathbf{\tilde{C}})\mathbf{x}\|^{2}\leq$ (33)
$\|\mathbf{H}(\mathbf{C})\mathbf{x}-\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}\|^{2}+\|\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}-\mathbf{H}(\mathbf{\tilde{C}})\mathbf{x}\|^{2}=$ (34)
$\alpha+\beta$ (35)
@@
E-A Covariance uncertainty error
The effect of covariance estimation errors on the stability of VNN is deterministic and, consequently, can be bounded by the square of the bound in Thm.˜1:
$\|\mathbf{H}(\mathbf{C})\mathbf{x}-\mathbf{H}(\mathbf{\hat{C}})\mathbf{x}\|^{2}\leq\frac{P^{2}k_{\textnormal{max}}^{2}}{t^{1-2\epsilon}}\mathcal{O}\left(N+\frac{\|\mathbf{C}\|^{2}\log(Nt)}{k_{\textnormal{min}}^{2}t^{4\epsilon}}\right)$ (36)
with probability at least $1-t^{-2\epsilon}-2\kappa N/t$ for any $\epsilon\in(0,1/2]$, where all the terms are defined in Thm.˜1.
E-B Covariance sparsification error
@@
Consider a random covariance matrix $\mathbf{\tilde{C}}_{r}=\mathbf{\hat{C}}+\mathbf{E}_{r}$, where $\mathbf{\tilde{C}}_{r}$ is a copy of the sample covariance matrix $\mathbf{\hat{C}}$ whose elements $\hat{c}_{ij}$ are dropped independently with probability $1-p_{ij}$ and $\mathbf{E}_{r}$ is its distance from $\mathbf{\hat{C}}$. Under the conditions of Def.˜4, i.e., edge perturbations are undirected (i.e., $\mathbf{E}_{r}$ is symmetric) and no perturbations occur on the diagonal (i.e., $\mathbf{E}_{r}$ has zeros on the diagonal), it holds that
@@
Each entry of the random error matrix $\mathbf{E}_{r}$ can be represented as $\mathbf{E}_{r}]_{ij}=-\delta_{ij}c_{ij}$, where $\delta_{ij}$ is a Bernoulli variable that is one with probability $1-p_{ij}$ and zero with probability $p_{ij}$. By performing the matrix multiplication, we can express each element of $\mathbf{E}_{r}^{2}$ and, consequently, its expectation, as
@@
$\mathsf{trace}(\mathbb{E}[\mathbf{E}_{r}^{2}])=\sum_{i=1}^{N}\sum_{n=1}^{N}\hat{c}_{in}\hat{c}_{ni}\mathbb{E}[\delta_{in}\delta_{ni}]=\sum_{i=1}^{N}\sum_{n=1}^{N}\hat{c}_{in}^{2}(1-p_{in}).$ (40)
@@
$\mathbb{E}\left[\sum_{r=1}^{K}\mathsf{trace}\left(\sum_{k=r}^{K}\sum_{l=r}^{K}h_{k}h_{l}\mathbf{E}_{r}\mathbf{\hat{C}}^{k+l-2r}\mathbf{E}_{r}\mathbf{\hat{C}}^{r-1}\mathbf{x}\mathbf{x}^{\mathsf{T}}\mathbf{\hat{C}}^{r-1}\right)\right]\leq$ (41)
$NP^{2}\|\mathbf{x}\|^{2}Q$ (42)
with $Q$ defined in (37).
@@
Following [36, eq. (B.13)-(B.15)], we rewrite the term in (41) as
@@
which holds for any square matrix $\mathbf{A}$ and positive semi-definite matrix $\mathbf{B}$ [75]. We note that $\mathbb{E}[\mathbf{E}_{r}^{2}]$ is positive semi-definite since $\mathbf{E}_{r}^{2}$ is the square of a symmetric matrix. Therefore, we use (44) to write
@@
From [36, eq. (B.24)], we have that the term within the norm is bounded by the generalized Lipschitz coefficient of the filter, i.e.,
@@
Then, using Lemma˜3, we have that
@@
Substituting these two terms in (46) leads to
@@
Similarly to Lemma˜3, we can express each element of the expected value of the matrix product $\mathbf{E}_{r_{1}}\mathbf{E}_{r_{2}}$ as
@@
From Lemma˜3 we have that
@@
$[\mathbb{E}[\mathbf{E}_{r_{1}}^{2}\mathbf{E}_{r_{2}}]]_{ij}=\sum_{m=1}^{N}\mathbb{E}[[\mathbf{E}_{r_{1}}^{2}]_{im}\hat{c}_{mj}\delta_{r_{2},mj}]=$ (55)
$\sum_{m=1}^{N}\sum_{n=1}^{N}\hat{c}_{in}\hat{c}_{nm}\hat{c}_{mj}\mathbb{E}[\delta_{r_{1},in}\delta_{r_{1},nm}\delta_{r_{2},mj}].$ (56)
@@
$\mathbb{E}[\delta_{r_{1},in}\delta_{r_{1},nm}\delta_{r_{2},mj}]=$ (57)
$\begin{cases}(1-p_{in})(1-p_{nm})(1-p_{mj})&\text{if }i\neq m\\ (1-p_{in})(1-p_{mj})&\text{if }i=m\end{cases}$ (58)
@@
$\mathsf{trace}(\mathbb{E}[\mathbf{E}_{r_{1}}^{2}\mathbf{E}_{r_{2}}])=$ (59)
$\sum_{i=1}^{N}\sum_{m=1}^{N}\sum_{n=1}^{N}\hat{c}_{in}\hat{c}_{nm}\hat{c}_{mi}\mathbb{E}[\delta_{r_{1},in}\delta_{r_{1},nm}\delta_{r_{2},mi}].$ (60)
Therefore, the trace is a sum of terms of quadratic and cubic order in the probability value, i.e., $\mathcal{O}((1-p_{1})(1-p_{2}))$ or $\mathcal{O}((1-p_{1})(1-p_{2})(1-p_{3}))$, where $p_{1},p_{2},p_{3}$ are generic probabilities values and the quadratic terms dominate the behavior since 1p<11-p<1. So,
@@
Consider the sample covariance matrix $\mathbf{\hat{C}}$ and a random sparsified covariance $\mathbf{\tilde{C}}_{r}=\mathbf{\hat{C}}+\mathbf{E}_{r}$ as in Def.˜4. We have that
@@
Similarly to Lemma˜3, we can write the expected value of each element of $\mathbf{E}^{3}_{r}$ as
$[\mathbb{E}[\mathbf{E}^{3}_{r}]]_{ij}=\sum_{m=1}^{N}\mathbb{E}[[\mathbf{E}_{r}^{2}]_{im}\hat{c}_{mj}\delta_{mj}]=$ (63)
$\sum_{m=1}^{N}\sum_{n=1}^{N}\hat{c}_{in}\hat{c}_{nm}\hat{c}_{mj}\mathbb{E}[\delta_{in}\delta_{nm}\delta_{mj}]$ (64)
@@
$\mathbb{E}[\delta_{in}\delta_{nm}\delta_{mi}]=$ (66)
$\begin{cases}(1-p_{in})(1-p_{nm})(1-p_{mi})&\text{if }i\neq m\land i\neq n\\ (1-p_{in})(1-p_{mi})&\text{if }i=m\oplus i=n\\ (1-p_{in})&\text{if }i=m=n\end{cases}$ (67)
where $\oplus$ denotes the xor operator. Now we note that, according to Def.˜4, the matrix $\mathbf{E}_{r}$ has zeros on the diagonal or, equivalently, $1-p_{ii}=0\quad\forall i$. As a consequence, the terms where $i=m=n$ in (65) are all zeros and the trace only contains terms of quadratic and cubic order in the probability value, i.e., $\mathcal{O}((1-p_{1})(1-p_{2}))$ or $\mathcal{O}((1-p_{1})(1-p_{2})(1-p_{3}))$, where $p_{1},p_{2},p_{3}$ are generic probabilities values and the quadratic terms dominate the behavior since 1p<11-p<1. Therefore,
@@
[…]
@@
We represent a random covariance matrix as $\mathbf{\tilde{C}}_{r}=\mathbf{\hat{C}}+\mathbf{E}_{r}$, where $\mathbf{E}_{r}$ is a random matrix that contains the deviation from the true covariance. Following [36, eq. (B.2)-(B.6)], we express the first term in (70) as
[…]